Status Report

NASA Space Station On-Orbit Status 16 Dec 2003

By SpaceRef Editor
December 16, 2003
Filed under , , ,
NASA Space Station On-Orbit Status 16 Dec 2003
iss

All ISS systems continue to function nominally, except as noted previously or below.

FE Alexander Kaleri continued the R&R (removal & replacement) of the BTA heat exchanger of the Russian segment’s SKV-2 air conditioner in the Service Module (SM).   [Today’s tasks consisted of demating (later remating) of the BITS2-12 onboard telemetry system from/to the SKV-2, replacement of the BTA, connecting it to the compressor (KU) and condenser unit (BK), etc.  Not all scheduled tasks could be completed, and an additional 2-hr. timeslot has been scheduled tomorrow for the Flight Engineer for closeout ops and ground-commanded checkout.]

In the U.S. segment (USOS), the leak check of two valves of the Node starboard hatch to the Airlock (A/L) was successfully completed with A/L repressurization and hatch opening.   [Afterwards, CDR Mike Foale reconfigured the A/L flexible air ducting to IMV (intermodular ventilation) return, stowed the tools, retransferred A/L cargo items from their temporary stowage in the Node to the A/L and reinstalled its PCS laptop.  Some questions of Mike’s during the ops regarding the proper configuration of the depress air return feedthrough line were clarified satisfactorily.  During the 8-hr. leak monitoring between the reduced-pressure (to ~560 mmHg {from 760 mmHg regular}) A/L and the increased-pressure (to ~770 mmHg) station interior, a minimal pressure drop of 3 mmHg was observed, which ranks well below the value of 10 mmHg in 8 hrs. used as criterion for a leak.]

Mike Foale continued the alteration work on the FOOT LEMS (Foot/Ground Reaction Forces during Space Flight/ Lower Extremity Monitoring Suit) garment pants started last month.  [He located two LEMS kits (Bowersox & Thomas) and used the sewing kit to perform pant alteration and a cable swap on Thomas’ LEMS pants.  LEMS consists of black Lycra biking tights with 20 electrodes and shoes fitted with insoles that measure impact forces on the bottom of the feet.]
 
The FE performed the daily routine maintenance of the SOZh life support system and prepared the regular IMS (inventory management system) “delta” file for automatic export/import to update the database.   [Specific questions by TsUP regarding the IMS file update today concerned entries of “Sputnik-SM” ham radio hardware and “Pilot” experiment equipment in a soft container in the FGB.]

The crew worked out according to their regular daily physical exercise program of 2.5 hrs on the TVIS treadmill, RED exerciser, CEVIS cycle ergometer and VELO bike (with load trainer).  To keep a closer watch on TVIS/treadmill health, the crew was requested to conduct the hitherto monthly wire rope inspection procedure now on a weekly basis. 

Foale performed a “ghost” software reload of the EXPRESS rack 1 (ER1) laptop computer (ELC), to recover it from a problem identified by him some time ago.

At 10:55am EST, Mike held a teleconference with ground specialists on the MSG PFMI (Pore Formation & Mobility Investigation), to discuss the experiment’s science background.  Meanwhile, PFMI troubleshooting on the ground continues.

During Increment 7, some traces of propylene glycol were detected in a condensate water sample.  Experts now believe that the source may have been U.S. body bath pouches.  The crew was asked to make sure to use only Russian wet towels, which are wetted with a cleaning solution, or U.S. shampoo, which is alcohol free.

Elektron Update:  The O2 generator is currently off, after its primary and backup pumps again shut down.  TsUP specialists are expected to have further troubleshooting steps tomorrow morning (GMT).  Current plan is to run the Elektron on its lowest power level of 16 amps, to preserve the membrane in the air/liquid separator of the Liquid Unit (BZh), which will have to be replaced.  The BZh’s certified life is one year, and the present unit was installed in November 2002.  A spare unit is on board.  Along with refreshes with Progress 12P O2 (the next tomorrow, 12/17, to increase ppO2 by 5-7 mmHg), the current arrangement will be maintained as long as possible.

Flight Attitude Update:   The station continues flight in LVLH YVV attitude (local vertical/local horizontal; +y-axis in velocity vector, i.e., flying “sidewise”) until 12/26.  The solar Beta angle is approximately -72 deg, and its magnitude is increasing, to peak tomorrow (12/17) at about -75 deg.  At these high Betas, the ISS does not experience eclipse, i.e., is always in sunlight. The extended insolation period leads to elevated temperatures on components such as radiator IMCAs (Integrated Motor Controller Assemblies) and RBVMs (Radiator Beam Valve Modules).  The USOS solar arrays are in directed position, i.e. non-suntracking, to conserve BGA (Beta Gimbal Assembly) life, and angled such that orbital drag is minimized. 

Today’s CEO targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for some viewing angles, were Lagos, Nigeria (nadir pass over this sixth largest urban center in the world, with 14.7 million people.  Looking on the sandspit on the seaward side of the coastal lagoon [Nigeria’s “Padre Island”], and opposite on the mainland), Ganges smog, India (a late-season tropical storm, affecting India from the Bay of Bengal, is expected to start dispersing the region-wide smog blanket that has been in place for two weeks.  Suggesting oblique views looking left ahead of the storm down the Ganges plain where visibility remains low [1-2 miles in Delhi].  The dynamics of smog dispersion are of interest), Rome, Italy (general view left of track.  Opportunity for views of smoking Etna Volcano under track about 60 secs later), Phoenix, AZ (good near-nadir view of Phoenix: looking a touch right for the whole urban region), and Gardner reef, Pacific (looking a touch right for this small atoll).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 2:00pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered Off.  Vozdukh CO2 scrubber is On (in Manual Mode 5/3).  U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode).  TCCS (trace contaminant control subsystem) is operating.  MCA (major constituents analyzer) is off (in Life Extending Mode).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On; SKV-2 is Off.
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 25.6; ppO2 (mmHg) — 141.9; ppCO2 (mmHg) — 3.8.
  • SM Transfer Compartment:  Pressure (mmHg) — 761; temperature (deg C) — 19.2.
  • FGB Cabin:  Pressure (mmHg) — 756; temperature (deg C) — 22.0.
  • Node:  Pressure (mmHg) — 744.52; temperature (deg C) — 23.9 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 746.94; temperature (deg C) — 24.8; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 747.04; temperature (deg C) — 24.5; shell heater temp (deg C) — 23.9, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 24.9
  • PMA-2:  Shell heater temp (deg C) — 11.4

(n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Directed Position (non-suntracking, “night glider”/”sun slicer” drag reduction mode). 
  • SM batteries:  Battery #8, formerly known as #7, is still disconnected in slot #8 for troubleshooting, off-line; all other batteries (7) are in “Partial Charge” mode.  
  • FGB batteries:  Battery #5 is off (capacity restoration mode, ROM); all other batteries (5) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22).

Propulsion System:

  • Total propellant load available: 3680 kg (8113 lb) as of 12/11/03  [SM(755) + FGB(2573) + Progress M(352) + Progress M-1(0)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH  YVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, y-axis in velocity vector [yaw: -89.9 deg, pitch: -8.9 deg, roll: 1.8 deg]), with CMG/TA (thruster assist) Momentum Management.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.

Robotics:

  • SSRMS/Canadarm2 based at MBS PDGF #2/LEE A, with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 10:43am EST [= epoch]):

  • Mean altitude — 371.4 km
  • Apogee — 375.5 km
  • Perigee — 367.3 km
  • Period — 92.0 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.0006037
  • Orbits per 24-hr. day — 15.66
  • Mean altitude loss in last 24 hours — 56 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 28953
  • For more on ISS orbit and worldwide naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html

SpaceRef staff editor.